OpenStudy (anonymous):

if 3/x is equal to 8x/x what is x?

zepdrix (zepdrix):

\[\huge \frac{3}{x}=\frac{8x}{x}\]Simplify the right side by canceling out the x's.\[\huge \frac{3}{x}=8\]Multiply both sides by x.\[\huge 3=8x\]Then to solve for x, divide by sides by 8.

OpenStudy (anonymous):

OMG YOU ARE A LIFE SAVER!!!!!!!!

OpenStudy (anonymous):

thank you~